OpenStudy (anonymous):

Emma receives $7750 and invests it in an account that earns 4% interest compounded continuously. What is the total amount of her investment after 5 years?

jimthompson5910 (jim_thompson5910):

A=Pe^(rt) A=7750*e^(0.04*5) A=7750*e^(0.2) A=7750*1.22140275816017 A=9465.87137574132 A=9465.87 So the total amount of the investment after 5 years is $9,465.87